OpenStudy (anonymous):

multiply and combine like terms if necessary. 4a(5a-1)^2

Parth (parthkohli):

Exponents first.\[\rm 4a \cdot (5a - 1)^2 = 4a \cdot{\Large (} (5a)^2 - 2(5a)(1) +(1)^2{\Large)} = 4a\cdot (25a^2 -10a+1) \]

Parth (parthkohli):

Use the distributive property to continue. I hope you can do the rest. :)